Find the probability of drawing an ace or a spade from a deck of 52 cards in a single draw.
1
Expert's answer
2018-04-12T05:02:11-0400
P(ace) = 4/52
P(spade) = 13/52
P(ace of spades) = 1/52
Then full probability will be
P(ace or spade) = (4 + 13 - 1)/52 = 4/13